OpenStudy (anonymous):

Fully factorize 3y^2-27 The answer is 3(y+3)(y-3), but why? I'm lost.

OpenStudy (anonymous):

3y^2-27 = 3(y^2-9)

OpenStudy (anonymous):

=3( y^2-3^2) 3(y+3)(y-3)

OpenStudy (anonymous):

a^2-b^2 = (a+b) (a-b)

OpenStudy (anonymous):

Thanks, makes a lot more sense to me now! (: